OpenStudy (pamstevens):

what is the complex root of x^2-9x^2+20=0

OpenStudy (unklerhaukus):

apply the quadratic formula

OpenStudy (pamstevens):

is that ax^2+bx+c=0?

OpenStudy (unklerhaukus):

Wait, is the question supposed to be x^2 - 9x^2 + 20 = 0 or x^2 - 9x + 20 = 0 ?

OpenStudy (sshayer):

there is no complex root as disc.=b^2-4ac>0

OpenStudy (unklerhaukus):

If there are two terms of x^2, then you should simplify them

OpenStudy (pamstevens):

It is x^2-9x^2+20=0

OpenStudy (unklerhaukus):

simplify x^2 - 9x^2 i.e. (1-9)x^2

OpenStudy (pamstevens):

so -8x^2+20=0

OpenStudy (unklerhaukus):

yeah

OpenStudy (unklerhaukus):

then get the x term to the other side

OpenStudy (pamstevens):

20=8x^2

OpenStudy (unklerhaukus):

now divide by 8

OpenStudy (pamstevens):

2.5=x^2

OpenStudy (unklerhaukus):

yes, but express 2.5 as an inproper fraction

OpenStudy (unklerhaukus):

5/2 = x^2 now take the square roots of both sides

OpenStudy (solomonzelman):

I think if the question is asking for complex roots, then the question is actually \(\color{black}{\displaystyle x^2-9x+20=0 }\), because the way it is written right now, the roots are not complex.

OpenStudy (solomonzelman):

Well, they are complex, but not complex exclusively.

OpenStudy (unklerhaukus):

the solution are real either way, and real numbers are a type of complex number

OpenStudy (solomonzelman):

oh yes, right. 4.5^2>20.

OpenStudy (solomonzelman):

My completing the square thinking gone a little off. Sorry:)

OpenStudy (pamstevens):

im not sure how to divide the square root

OpenStudy (unklerhaukus):

x^2 = 5/2 x = ±√(5/2)

OpenStudy (unklerhaukus):

that's two roots x = √(5/2) and x = -√(5/2). both of these are real numbers

OpenStudy (pamstevens):

so then that is the final answers correct? One positive and the negative?

OpenStudy (unklerhaukus):

yeah, they don't simplify into nicer forms unfortunately

OpenStudy (pamstevens):

Thank you I appreciate your help.